OpenStudy (anonymous):

f(s) = (s^3-1) / (s^2-1) How do you divide something like this? Does it factor? If so, how?

OpenStudy (anonymous):

yes it factors

OpenStudy (anonymous):

difference of cubes/difference of squares

OpenStudy (anonymous):

I'm trying to give you a hint

OpenStudy (anonymous):

Well so i understand that it ends up being 3/2 but can you show me the steps involved? Because all i get so far is that the denominator can factor to (s+1)(s-1) but what can the numerator factor to?

OpenStudy (anonymous):

Everybody wants to help :)

OpenStudy (anonymous):

nuemerato = (s-1)(s^2 +s + 1)

OpenStudy (anonymous):

- if my memory serves me right!!

OpenStudy (anonymous):

Yes try to factor: f(s) = (s^3 -1)/(s^2 -1) f(s) = (s -1) (s^2 +s +1) /[(s-1)(s+1)] <--cross out s-1 f(s)= (s^2 +s +1)/ (s +1) and im not sure how you end up with 3/2 though :D

OpenStudy (anonymous):

I think he's thinking about taking the limit with coefficients in the numerator and denominator with polynomials of the same degree.

OpenStudy (anonymous):

you can go further to break it up to f(x)= (s^2 +s )/ (s+1) + 1/(s+1) f(x) = s(s+1)/(s+1) + 1/(s+1) f(x) = s + 1/(s+1)

OpenStudy (anonymous):

if s =1 then it is 3/2

OpenStudy (anonymous):

yeah the limit is 3/2 when x->1 that's what i meant.... but i'm not seeing how (s-1)(s^2 + s +1) = s^3 - 1

OpenStudy (anonymous):

oh, so this is a limit question xD

OpenStudy (anonymous):

I knew it

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

\[\color{red}{\frac{s^3-1}{s^2-1}=s + \frac{1}{s-1}}\text{ right?}\]

OpenStudy (amistre64):

mighta got a - and + switched

OpenStudy (anonymous):

It does?

OpenStudy (anonymous):

bovice: that's just a property of difference of cube... they have formula for it (a^3 - b^3) = (a - b) (a^2 +ab +b^2)

OpenStudy (anonymous):

f(x) = X^3 -1 then f(1) = 0 so x-1 is a factor

OpenStudy (anonymous):

divide x-1 into x^3 -1 gives x^2 + x+ 1

OpenStudy (anonymous):

yeah (s-1) will be cross out instead of x+1

OpenStudy (anonymous):

s +1 i meant

OpenStudy (anonymous):

I guess so...... still confused on this but i dont really know the steps for diving polynomials

OpenStudy (amistre64):

dividing polys are the same as dividing numbers; if you recall how to divide longhand

OpenStudy (anonymous):

I knew he would be confused...too many people giving their way of dividing them

OpenStudy (anonymous):

hmmm.... x-1 / x^3 -1 x goes into x^3 x^2 times....

OpenStudy (anonymous):

that's as far as i can get....dunno what to do next

OpenStudy (anonymous):

the one thng you have to watch when u divide polys is the powers must go down by 1 so when dividing into x^3 - 1 u need to write it as x^3 + 0x^2 + 0x - 1

OpenStudy (anonymous):

amistre64 would use an online classroom to show you step by step but he needs an instruction manual to use the whiteboard, lol

OpenStudy (amistre64):

\(\color{blue}{\text{s}}\) \(\color{green}{\text{------------}}\) \(\color{green}{\text{s^2 - 1| s^3 - 1}}\) \(\color{red}{\text{-s^3 +s}}\) \(\color{black}{\text{----------}}\) \(\color{green}{\text{s - 1}}\) <- remainder

OpenStudy (amistre64):

whiteboards are for sissys :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Wow, how creative

OpenStudy (mathteacher1729):

Here is a nice summary of the difference of cubes and difference of squares. :) http://www.wtamu.edu/academic/anns/mps/math/mathlab/int_algebra/int_alg_tut29_specfact.htm

OpenStudy (anonymous):

amistre64 proves he doesn't need a whiteboard

OpenStudy (amistre64):

real mathmatickers \(\color{orange}{\textbf{ do it in color}}\) :)

OpenStudy (anonymous):

I don't know amistre64....Latex wastes manhours

OpenStudy (anonymous):

When you waste manhours, you become a sissy

OpenStudy (amistre64):

lol , so true :)

OpenStudy (anonymous):

alright i found out how to using that link....thanks

OpenStudy (anonymous):

So yeah, should be \[s + {1 \over s+1}\] when all is said and done.

OpenStudy (amistre64):

that looked better in preview :)

OpenStudy (anonymous):

He used over instead of frac

myininaya (myininaya):

but everyone's favorite thing to do is use l'hospital's rule

OpenStudy (anonymous):

You're so right. I love l'hopital rule

OpenStudy (anonymous):

Wait, you spelled hospital wrong...

myininaya (myininaya):

lol its the thought that counts

OpenStudy (anonymous):

l'hopitals rule why didnt i think of that.... x^3 -1 = 3x^2 and x^2-1 = 2x so 3(1^2) / 2(1) = 3/2 hahahaha woooo

myininaya (myininaya):

lol

OpenStudy (anonymous):

Because you can only use it when you have 0/0 or infinity/infinity

myininaya (myininaya):

we do have 0/0

OpenStudy (anonymous):

lol, you're right...nobody thought to do it. I blame the OP

OpenStudy (anonymous):

Well he originally didn't say he was looking for a limit.

myininaya (myininaya):

right im glad i came late in the conversation

myininaya (myininaya):

:)

OpenStudy (anonymous):

exactly...but he did mention it though, but by the time he did, everyone was already working on the hard way...

OpenStudy (anonymous):

Hey Bovice, I have your alternate solution. factor the numerator and the denominator as before but then substitute s = 1 into the function to get 3/2.

OpenStudy (anonymous):

lim x --> 1 of ((s-1)(s^2 + s+1))/((s-1)(s+1)

OpenStudy (anonymous):

Simplify: lim x --> 1 (s^2+s+1)/(s+1)

OpenStudy (anonymous):

substitute 1: (1+1+1)/(1+1) = 3/2